OpenStudy (luigi0210):

Use cylindrical shells:

OpenStudy (anonymous):

You don't really want to use shells for this.

OpenStudy (luigi0210):

Sadly, I have to.. it's method they told us to use.

OpenStudy (anonymous):

\[\begin{array}{rcl} 1<&y&<2 \\ -\sqrt{2-y}<&x&<\sqrt{2-y} \end{array} \]

OpenStudy (anonymous):

\You have to use \[ 2\pi \int rh~dy \]

OpenStudy (anonymous):

In this case \(r = |y-1|\), however this will always be positive so it simplifies to \(r=y-1\)

OpenStudy (anonymous):

And \(h= \sqrt{2-y} - (-\sqrt{2-y}) = 2\sqrt{2-y}\)

OpenStudy (anonymous):



OpenStudy (anonymous):

To get \(h\), I used \[ y= 2-x^2 \implies x^2=2-y \implies x=\pm \sqrt{2-y} \]

OpenStudy (anonymous):

All together: \[ V = 2\pi \int_1^2(y-1)(2\sqrt{y-2})~dy \]

OpenStudy (luigi0210):

Alright, just needed to check the set up, thanks wio :)